# Coordinate In and Value Out: Training Flow Transformers in Ambient Space

### v1

@@ Abstract
Flow matching models have emerged as a powerful method for generative modeling on domains like images or videos, and even on unstructured data like 3D point clouds. These models are commonly trained in two stages: first, a data compressor (i.e. a variational auto-encoder) is trained, and in a subsequent training stage a flow matching generative model is trained in the low-dimensional latent space of the data compressor. This two stage paradigm adds complexity to the overall training recipe and sets obstacles for unifying models across data domains, as specific data compressors are used for different data modalities. To this end, we introduce Ambient Space Flow Transformers (ASFT), a domain-agnostic approach to learn flow matching transformers in ambient space, sidestepping the requirement of training compressors and simplifying the training process. We introduce a conditionally independent point-wise training objective that enables ASFT to make predictions continuously in coordinate space. Our empirical results demonstrate that using general purpose transformer blocks, ASFT effectively handles different data modalities such as images and 3D point clouds, achieving strong performance in both domains and outperforming comparable approaches. ASFT is a promising step towards domain-agnostic flow matching generative models that can be trivially adopted in different data domains.
@@ 1 Introduction
Recent advances in generative modeling have enabled learning complex data distributions by combining both powerful architectures and training objectives. In particular, state-of-the-art approaches for image  (Esser et al., 2024), video (Dai et al., 2023) or 3D point cloud  (Vahdat et al., 2022) generation are based on the concept of iteratively transforming data into Gaussian noise. Diffusion models were originally proposed following this idea and pushing the quality of generated samples in many different domains, including images (Dai et al., 2023; Rombach et al., 2022), 3D point clouds (Luo & Hu, 2021), graphs (Hoogeboom et al., 2022) and video (Ho et al., 2022a). More recently, flow matching (Lipman et al., 2023) and stochastic interpolants (Ma et al., 2024) have been proposed as generalized formulations of the noising process, moving from stochastic gaussian diffusion processes to general paths connecting a base (e.g. Gaussian) and a target (e.g. data) distribution.
In practice, these iterative refinement approaches are commonly applied in a low-dimensional latent space obtained from a pre-trained compressor model. Therefore, the training process for these approaches is composed of two independent training stages: in the first stage, a compressor (e.g. VAE (Vahdat et al., 2022), VQVAE (Ramesh et al., 2022), VQGAN (Rombach et al., 2022)) model is trained, using architectures that are specific to the data domain (i.e. ConvNets for image data (Rombach et al., 2022), PointNet for point clouds (Vahdat et al., 2022), etc. ) enforcing a bottleneck on the data dimensionality, with the goal of reducing compute cost of training the subsequent stage. In the second stage, general purpose transformer architectures are used for the generative modeling step (Peebles & Xie, 2023; Ma et al., 2024; Esser et al., 2024), where the distribution of latents is learnt. This type of generative modeling in latent space has become popular in the community due to its computational efficiency benefits obtained from compressed data dimensionality.
However, latent space generative modeling is not without drawbacks. An obvious shortcoming is that latent space generative models cannot benefit from end-to-end optimization, as data compressors and the downstream generative models are trained separately. In particular, two stage approaches are more complex to implement than single stage models and involve tuning several hyper-parameters that can have a big impact on final performance (Rombach et al., 2022): spatial reduction ratio, adversarial loss weights or KL terms in VAEs (Rombach et al., 2022) . As an illustrative example, setting a high KL weight makes the problem of learning a distribution of latents trivial, yet results in very poor generation results. A very small KL weight on the other hand allows for great reconstruction performance for the first stage but fails to induce a suitable latent space for generative modeling (e.g. a dirac delta for each training sample in latent space). Our goal in this paper is to provide a single training stage approach that is domain-agnostic and simple to implement in practice, thus dispensing with the complexities of two stage training recipes and enabling modeling of different data modalities in ambient (i.e. data) space.
It is worth noting that training diffusion or flow matching models in ambient space is indeed possible when using domain specific architecture designs and training recipes. In the image domain, approaches have exploited its dense nature and applied cascaded U-Nets Ho et al. (2021; 2022b), joint training of U-Nets at multiple resolutions Gu et al. (2023), multi-scale losses (Hoogeboom et al., 2023) or U-Net transformer hybrids architectures (Crowson et al., 2024), obtaining strong results. However, developing strong domain-agnostic models, using general purposes architectures that can be applied across different data domains remains an important open problem.
@@
Figure 1: (a) High level overview of ASFT using the image domain as an example. Our model can be interpreted as an encoder-decoder model where the decoder makes predictions independently for each coordinate-value pair given ${\bm{z}}_{f_{t}}$. For 3D point clouds, the coordinate and value are equivalent and their dimensions change, but the model is the same. (b) Images generated by ASFT trained on ImageNet-256 and (c) 3D point clouds ($32k$ points) generated by training ASFT on Objaverse.
@@
In this paper, we answer a three part question: Can we learn flow matching models in ambient space, in a single training stage and using a domain agnostic architecture? Our goal is to unify different data domains under the same training recipe. To achieve this, we introduce Ambient Space Flow Transformers (ASFT), see Fig. 1(a). ASFT makes progress towards the goal of unifying flow matching generative modeling across data domains. The key component of our approach is a conditionally independent point-wise training objective that enables training in ambient space and can be densely (e.g. continuously) evaluated during inference. In the image domain, this means that we model the probability of a pixel value given its coordinate, which provides granular and precise control over image synthesis, allowing to generate images at different resolution than the one used during training (see Fig. 4(a)). We show ImageNet-256 generated samples from ASFT in Fig. 1(b) and 3D point clouds from Objaverse in Fig. 1(c) (see additional samples in Fig. 6, 7, 8, 9, 10). Our contributions are summarized as follows:
@@
  • We propose ASFT, a flow matching generative transformer that works on ambient space to enable single stage generative modeling on different data domains.
@@
  • Our results show that ASFT, though domain-agnostic, achieves competitive performance on image and 3D point cloud generation compared with strong domain-specific baselines.
@@ 2 Related Work
Diffusion models have been the major catalyzer of progress in generative modeling, these approaches learn to reverse a forward process that gradually adds Gaussian noise to corrupt data samples (Ho et al., 2020). Diffusion models are notable for their simple and robust training objective. Extensive research has explored various formulations of the forward and backward processes (Song et al., 2021a; Rissanen et al., 2022; Bansal et al., 2022), particularly in the image domain. In addition, different denoising networks have been proposed for different data domains like images (Nichol & Dhariwal, 2021), videos (Ho et al., 2022a), and geometric data (Luo & Hu, 2021). More recently, flow matching (Liu et al., 2023; Lipman et al., 2023) and stochastic interpolants (Ma et al., 2024) have emerged as flexible formulations that generalized Gaussian diffusion paths, allowing to define different paths to connect a base and a target distribution. These types of models have shown incredible results in the image domain (Ma et al., 2024; Esser et al., 2024) when coupled with transformer architectures (Vaswani et al., 2017) to model distributions in latent space learnt by data compressors (Peebles & Xie, 2023; Ma et al., 2024; Rombach et al., 2022; Vahdat et al., 2022; Zheng et al., 2023; Gao et al., 2023). Note that these approaches train two separate stages/models: first training the data compressor (e.g. VAE (Vahdat et al., 2022), VQVAE (Ramesh et al., 2022), VQGAN (Rombach et al., 2022)) and then the generative model, requiring careful hyper-parameter tuning.
In an attempt to unify generative modeling across various data domains, continuous data representations [^1] These are also often referred to as implicit neural representation, neural maps or neural operators have shown potential in different approaches: From Data to Functa (Functa) (Dupont et al., 2022a), Generative Manifold Learning (GEM) (Du et al., 2021a), and Generative Adversarial Stochastic Process (GASP) (Dupont et al., 2022b) have studied the problem of generating continuous representations of data. More recently Infinite Diffusion (Bond-Taylor & Willcocks, 2023) and PolyINR (Singh et al., 2023) have shown great results in the image domain by modeling images as continuous functions. However, both of these approaches make strong assumptions about image data. In particular, (Bond-Taylor & Willcocks, 2023) interpolates sparse pixels to an euclidean grid to then process it with a U-Net. On the other hand, (Singh et al., 2023) uses a patching and 2D convolution in the discriminator. Our approach also relates to DPF Zhuang et al. (2023), a diffusion model that acts on function coordinates and can be applied in different data domains on a grid at low resolutions (i.e. 64$\times$64). Our approach is able to deal with higher resolution functions (e.g. 256x256 vs. 64x64 resolution images) on large scale datasets like ImageNet, while also tackling unstructured data domains that do not live on an Euclidean grid (e.g. like 3D point clouds).
@@ 3 Method
3.1 Data as Coordinate $\rightarrow$ Value Maps
@@
In practice, analytical forms for these maps $f$ are unknown. In addition, different from previous approaches (Dupont et al., 2022a; Du et al., 2021a), we do not assume that parametric forms of these maps can be obtained, since that would involve a separate training stage fitting an MLP to each map (Dupont et al., 2022a; Bauer et al., 2023; Du et al., 2021a). As a result, we assume we are only given sets of corresponding coordinate and value pairs resulting from observing these maps at a particular sampling rate (e.g. at a particular resolution in the image case). Therefore, we need a to develop an end-to-end approach that can take these collections of coordinate-value sets as training data.
@@ 3.2 Flow Matching and Stochastic Interpolants
${\bm{u}}_{t}(f)=\mathbb{E}[d_{t}{f_{t}}|f_{t}=f]=d_{t}\alpha_{t}\mathbb{E}[f_{0}|f_{t}=f]+d_{t}\sigma_{t}\mathbb{E}[\epsilon|f_{t}=f].$ (3)
@@
3.3 Flow Matching for Coordinate-Value Sets
We now turn to the task of formulating a flow matching training objective for data distributions of maps $f$. We recall that in practice we do not have access to an analytical or parametric form for these maps $f$, and we are only given sets of corresponding coordinate ${\bm{x}}_{f}$ and value ${\bm{y}}_{f}$ pairs resulting from observing the mapping at a particular rate. As a result, we need to formulate a training objective that can take these sets of coordinate-value as training data.
@@ 3.3 Flow Matching for Coordinate-Value Sets
$L_{\text{CICFM}}=\mathbb{E}_{t\sim\mathcal{U}[0,1],f\sim q(f),\epsilon\sim\mathcal{N}(0,{\mathbf{I}})}||{\bm{v}}_{\theta}({\bm{x}}_{f_{t}},{\bm{y}}_{f_{t}},t|{\bm{z}}_{f_{t}})-{\bm{u}}_{t}({\bm{x}}_{f},{\bm{y}}_{f}|\epsilon)||_{2}^{2},$ (4)
where the target velocity field ${\bm{u}}_{t}({\bm{x}},{\bm{y}}|\epsilon)$ is defined as a rectified flow (Liu et al., 2023; Lipman et al., 2023) or linear interpolant path (Ma et al., 2024):
${\bm{u}}_{t}({\bm{x}}_{f},{\bm{y}}_{f}|\epsilon)=(1-t)\epsilon+t{\bm{y}}_{f}.$ (5)
@@ 3.4 Network Architecture
Figure 2: Architecture of our proposed ASFT for different data domains including images and 3D point clouds. Note that models are trained for each data domain separately. Each spatial aware latent takes in a subset of neighboring context coordinate-value sets in coordinate space. The latents are then updated through self-attention. Decoded coordinate-value pairs cross attend to the updated latents ${\bm{z}}_{f_{t}}$ to decode the corresponding velocity.
We base our model on the general PerceiverIO design (Jaegle et al., 2022) which provides a flexible architecture to handle coordinate-value sets of large cardinality (i.e. large number of pixels in an image). Fig. 2 illustrates the architectural pipeline of ASFT. At a high level, our encoder network takes a set of coordinate-value pairs and encodes them to learnable latents through cross-attention. These latents are then updated through several self-attention blocks to provide the final latents ${\bm{z}}_{f_{t}}\in\mathbb{R}^{L\times D}$ To decode the velocity field for a given coordinate-value pair we perform cross attention to ${\bm{z}}_{f_{t}}$, generating the final point-wise prediction for the velocity field ${\bm{v}}_{\theta}({\bm{x}}_{f_{t}},{\bm{y}}_{f_{t}},t|{\bm{z}}_{f_{t}})$.
The encoder of a vanilla PerceiverIO relies solely on cross-attention to the latents $z_{f_{t}}\in\mathbb{R}^{L\times D}$ to learn spatial connectivity patterns between input and output elements, which we found to introduce a strong bottleneck during training. To ameliorate this, we make two key modifications to boost the performance. Firstly, our encoder utilizes spatial aware latents where each latent is assigned a “pseudo” coordinate. Coordinate-value pairs are assigned to different latents based on their distances on coordinate space. During encoding, coordinate-value pairs interact with their assigned latents through cross-attention. In particular, the learnable latent $z_{f_{t}}$ cross-attends to input coordinate-value pairs of noisy data at a given timestep $t$. Latent vectors are spatial-aware, this means that each of the $L$ latents only attends to a set of neighboring coordinate-value pairs. Latent vectors are then updated using several self-attention blocks. These changes in the encoder allow the model to effectively utilize spatial information while also saving compute when encoding large coordinate-value sets on ambient space. In the decoder, a given coordinate-value pair cross attends to ${\bm{z}}_{f_{t}}$ as in the original PerceiverIO. However, we found that a multi-level decoding strategy, which not only cross attends to the latents in the final layer but also the latent from the intermediate self-attentions layer is helpful. In particular, a given coordinate-value pair cross attends to latents from subsequent encoder layers sequentially to progressively refine the prediction. Finally, following previous work (Peebles & Xie, 2023; Ma et al., 2024), we apply AdaLN-Zero blocks for conditioning both on timestep $t$ and class labels whenever needed (e.g. for ImageNet experiments). More architectural details can be found in App. A.

### v2

# INRFlow: Flow Matching for INRs in Ambient Space

@@ Abstract
Flow matching models have emerged as a powerful method for generative modeling on domains like images or videos, and even on irregular or unstructured data like 3D point clouds or even protein structures. These models are commonly trained in two stages: first, a data compressor is trained, and in a subsequent training stage a flow matching generative model is trained in the latent space of the data compressor. This two-stage paradigm sets obstacles for unifying models across data domains, as hand-crafted compressors architectures are used for different data modalities. To this end, we introduce INRFlow, a domain-agnostic approach to learn flow matching transformers directly in ambient space. Drawing inspiration from INRs, we introduce a conditionally independent point-wise training objective that enables INRFlow to make predictions continuously in coordinate space. Our empirical results demonstrate that INRFlow effectively handles different data modalities such as images, 3D point clouds and protein structure data, achieving strong performance in different domains and outperforming comparable approaches. INRFlow is a promising step towards domain-agnostic flow matching generative models that can be trivially adopted in different data domains.
Machine Learning, ICML
@@ 1 Introduction
Recent advances in generative modeling have enabled learning complex data distributions by combining both powerful architectures and training objectives. In particular, state-of-the-art approaches for image  (Esser et al., 2024), video (Dai et al., 2023) or 3D point cloud  (Vahdat et al., 2022) generation are based on the concept of iteratively transforming data into Gaussian noise. Diffusion models were originally proposed following this idea and pushing the quality of generated samples in many different domains, including images (Dai et al., 2023; Rombach et al., 2022; Preechakul et al., 2022), 3D point clouds (Luo & Hu, 2021), graphs (Hoogeboom et al., 2022) and video (Ho et al., 2022a). More recently, flow matching (Lipman et al., 2023) and stochastic interpolants (Ma et al., 2024) have been proposed as generalized formulations of the noising process, moving from stochastic gaussian diffusion processes to general paths connecting a base (e.g. Gaussian) and a target (e.g. data) distribution.
@@
Figure 1: (a) High level overview of INRFlow using the image domain as an example. Our model can be interpreted as an encoder-decoder model where the decoder makes predictions independently for each coordinate-value pair given ${\bm{z}}_{f_{t}}$. For different data domains, the coordinate and value dimensionality changes, but the model is kept the same. (b) Samples generated by INRFlow trained on ImageNet 256$\times$256. (c) Image-to-3D point clouds generated by training INRFlow on Objaverse (Deitke et al., 2023). (d) Protein structures generated by INRFlow trained on SwissProt (Boeckmann et al., 2003). GT protein structures are depicted in green while the generated structures by INRFlow are show in orange.
In practice, these iterative refinement approaches are commonly applied in a latent space obtained from a pre-trained compressor model. Thus, the training recipe consists of two independent training stages: compressor (auto-encoder) training and subsequent generative modeling in latent space. General purpose transformer architectures have been used in the generative modeling step in latent space (Peebles & Xie, 2023; Ma et al., 2024; Esser et al., 2024). However, the first stage compressor uses architectures that are specific to the data domain, requiring hand-crafted inductive biases (i.e. ConvNets for image data (Rombach et al., 2022), PointNet for point clouds (Vahdat et al., 2022), Evoformer for protein structures (Jumper et al., 2021)).
We see this as one of the core issues preventing the ML community to develop truly domain-agnostic generative models that can be applied to different data domains in a trivial manner. Our goal in this paper is to provide a powerful single training stage approach that is domain-agnostic and simple to implement in practice, thus dispensing with the complexities of two-stage training recipes and enabling modeling of different data modalities directly in ambient (i.e. data) space.
It is worth noting that training diffusion or flow matching models in ambient space is indeed possible when using domain specific architecture designs and training recipes. In the image domain, approaches have exploited its dense nature and applied cascaded U-Nets (Ho et al., 2021, 2022b), joint training of U-Nets at multiple resolutions (Gu et al., 2023), multi-scale losses (Hoogeboom et al., 2023) or U-Net transformer hybrids architectures (Crowson et al., 2024), obtaining strong results. However, developing strong domain-agnostic models, using general purposes architectures that can be applied across different data domains remains an important open problem.
@@
In this paper, we answer a three part question: Can we learn flow matching models in ambient space, in a single training stage and using a domain agnostic architecture? Our goal is to unify different data domains under the same training recipe. To achieve this, we introduce INRFlow (INRFlow), see Fig. 1(a). INRFlow makes progress towards the goal of unifying flow matching generative modeling across data domains. Drawing inspiration from INRs, we formulate a conditionally independent point-wise training objective that enables training directly in ambient space and can be densely (e.g. continuously) evaluated during inference. In the image domain, this means that INRFlow models the probability of a pixel value given its coordinate (i.e. the probabilistic extension of an INR), allowing to generate images at different resolution than the one used during training (see Fig. 4(a)). We show generated samples from INRFlow trained on ImageNet-256 in Fig. 1(b), image-to-3D on Objaverse in Fig. 1(c) and protein structures on SwissProt Fig. 1(d) (see additional samples in Fig. 9, 12, 13). Our contributions are summarized as follows:
@@
  • We propose INRFlow, a flow matching generative transformer that works on ambient space to enable single stage generative modeling on different data domains.
@@
  • Our results show that INRFlow, though domain-agnostic, achieves competitive performance on image and 3D point cloud generation compared with strong domain-specific baselines.
@@ 2 Related Work
Diffusion models have been the major catalyzer of progress in generative modeling, these approaches learn to reverse a forward process that gradually adds Gaussian noise to corrupt data samples (Ho et al., 2020). Diffusion models are notable for their simple and robust training objective. Extensive research has explored various formulations of the forward and backward processes (Song et al., 2021; Rissanen et al., 2022; Bansal et al., 2022), particularly in the image domain. In addition, different denoising networks have been proposed for different data domains like images (Nichol & Dhariwal, 2021), videos (Ho et al., 2022a), and geometric data (Luo & Hu, 2021). More recently, flow matching (Liu et al., 2023; Lipman et al., 2023) and stochastic interpolants (Ma et al., 2024) have emerged as flexible formulations that generalized Gaussian diffusion paths, allowing to define different paths to connect a base and a target distribution. These types of models have shown incredible results in the image domain (Ma et al., 2024; Esser et al., 2024) when coupled with transformer architectures (Vaswani et al., 2017) to model distributions in latent space learnt by data compressors (Peebles & Xie, 2023; Ma et al., 2024; Rombach et al., 2022; Vahdat et al., 2022; Zheng et al., 2023; Gao et al., 2023). Note that these data compressors use domain specific architectures with hand-crafted inductive biases.
In an attempt to unify generative modeling across various data domains, continuous data representations (also referred to as implicit neural representation, neural fields or neural operators) have shown potential in different approaches: From Data to Functa (Functa) (Dupont et al., 2022a), Generative Manifold Learning (GEM) (Du et al., 2021a), and Generative Adversarial Stochastic Process (GASP) (Dupont et al., 2022b) have studied the problem of generating continuous representations of data. More recently Infinite Diffusion (Bond-Taylor & Willcocks, 2023) and PolyINR (Singh et al., 2023) have shown great results in the image domain by modeling images as continuous functions. However, both of these approaches make strong assumptions about image data. In particular, (Bond-Taylor & Willcocks, 2023) interpolates sparse pixels to an euclidean grid to then process it with a U-Net. On the other hand, (Singh et al., 2023) uses a patching and 2D convolution in the discriminator. Our approach also relates to DPF (Zhuang et al., 2023), a diffusion model that acts on function coordinates and can be applied in different data domains on a grid at low resolutions (i.e. 64$\times$64). Our approach is able to deal with higher resolution functions (e.g. 256x256 vs. 64x64 resolution images) on large scale datasets like ImageNet, while also tackling unstructured data domains that do not live on an Euclidean grid (e.g. like 3D point clouds and protein structures).
@@ 3 Method
3.1 Data as Continuous Coordinate $\rightarrow$ Value Maps
@@
In practice, analytical forms for these maps $f$ are unknown. In addition, different from previous approaches (Dupont et al., 2022a; Du et al., 2021a), we do not fit separate INRs to each data sample via reconstruction, since that would involve a separate training stage fitting an MLP for each map (Dupont et al., 2022a; Bauer et al., 2023; Du et al., 2021a). As a result, we assume we are only given sets of corresponding coordinate and value pairs resulting from observing these maps at a particular sampling rate (e.g. at a particular resolution in the image case). In the following, we develop an end-to-end approach that can directly take these coordinate-value sets as training data and train a model that extends INRs to the probabilistic setting.
@@ 3.2 Flow Matching and Stochastic Interpolants
${\bm{u}}_{t}(f)=\mathbb{E}[d_{t}{f_{t}}|f_{t}=f]=$
$=d_{t}\alpha_{t}\mathbb{E}[f_{0}|f_{t}=f]+d_{t}\sigma_{t}\mathbb{E}[\epsilon|f_{t}=f].$ (3)
@@
3.3 INRFlow
We now turn to the task of formulating a flow matching training objective for data distributions of maps $f$. We recall that in practice we do not have access to an analytical or parametric form for these maps $f$ (e.g. we do not requite to pretrain INRs for each training sample), and we are only given sets of corresponding coordinate ${\bm{x}}_{f}$ and value ${\bm{y}}_{f}$ pairs resulting from observing the mapping at a particular rate. As a result, we need to formulate a training objective that can take these sets of coordinate-value as training data.
@@ 3.3 INRFlow
$\mathbb{E}_{t\sim\mathcal{U}[0,1],f\sim q(f)}||{\bm{v}}_{\theta}({\bm{x}}_{f_{t}},{\bm{y}}_{f_{t}},t|{\bm{z}}_{f_{t}})-{\bm{u}}_{t}({\bm{x}}_{f},{\bm{y}}_{f}|\epsilon)||_{2}^{2},$ (4)
where the target velocity field ${\bm{u}}_{t}({\bm{x}},{\bm{y}}|\epsilon)$ is defined as a rectified flow (Liu et al., 2023; Lipman et al., 2023; Ma et al., 2024): ${\bm{u}}_{t}({\bm{x}}_{f},{\bm{y}}_{f}|\epsilon)=\frac{\epsilon-{\bm{y}}_{f_{t}}}{1-t}$.
@@ 3.4 Network Architecture
Figure 2: Architecture of our proposed INRFlow for different data domains including images and 3D point clouds. Note that models are trained for each data domain separately. Each spatial aware latent takes in a subset of neighboring context coordinate-value sets in coordinate space. The latents are then updated through self-attention. Decoded coordinate-value pairs cross attend to the updated latents ${\bm{z}}_{f_{t}}$ to decode the corresponding velocity.
We base our model on the general PerceiverIO design (Jaegle et al., 2022), Fig. 2 illustrates the architectural pipeline of INRFlow. At a high level, our encoder network takes a set of coordinate-value pairs and encodes them to learnable latents through cross-attention. These latents are then updated through several self-attention blocks to provide the final latents ${\bm{z}}_{f_{t}}\in\mathbb{R}^{L\times D}$ To decode the velocity field for a given coordinate-value pair we perform cross attention to ${\bm{z}}_{f_{t}}$, generating the final point-wise prediction for the velocity field ${\bm{v}}_{\theta}({\bm{x}}_{f_{t}},{\bm{y}}_{f_{t}},t|{\bm{z}}_{f_{t}})$.
The encoder of a vanilla PerceiverIO relies solely on cross-attention to the latents $z_{f_{t}}\in\mathbb{R}^{L\times D}$ to learn spatial connectivity patterns between input and output elements, which we found to introduce a strong bottleneck during training. To ameliorate this, we make a key modifications to boost the performance. Firstly, our encoder utilizes spatial aware latents where each latent is assigned a “pseudo” coordinate. Coordinate-value pairs are assigned to latents based on their distances on coordinate space. During encoding, coordinate-value pairs interact with their assigned latents through cross-attention, this means that each of the $L$ latents only attends to a set of neighboring coordinate-value pairs. Latent vectors are then updated using several self-attention blocks. These changes in the encoder allow the model to effectively utilize spatial information while also saving compute when encoding large coordinate-value sets on ambient space.